Becton Labs, Inc., produces various chemical compounds for industrial use. One compound, called Fludex, is prepared using an elaborate distilling process. The company has developed standard costs for one unit of Fludex, as follows:

Standard Quantity Standard Price
or Rate
Standard Cost
  Direct materials    2.20 ounces $ 25.00 per ounce $ 55.00   
  Direct labor    0.50 hours $ 15.00 per hour    7.50   
  Variable manufacturing overhead    0.50 hours $ 3.00 per hour    1.50   
    
$ 64.00   
    
    During November, the following activity was recorded relative to production of Fludex:

   

a. Materials purchased, 12,000 ounces at a cost of $282,000.
b.

There was no beginning inventory of materials; however, at the end of the month, 2,750 ounces of material remained in ending inventory.

c.

The company employs 25 lab technicians to work on the production of Fludex. During November, they worked an average of 110 hours at an average rate of $11.50 per hour.

d.

Variable manufacturing overhead is assigned to Fludex on the basis of direct labor-hours. Variable manufacturing overhead costs during November totaled $2,400.

e. During November, 4,100 good units of Fludex were produced .
Required:
1. For direct materials:
a.

Compute the price and quantity variances.

b. The materials were purchased from a new supplier who is anxious to enter into a long-term purchase contract. Would you recommend that the company sign the contract?
2. For direct labor:
a.

Compute the rate and efficiency variances

b.

In the past, the 25 technicians employed in the production of Fludex consisted of 5 senior technicians and 20 assistants. During November, the company experimented with fewer senior technicians and more assistants in order to save costs. Would you recommend that the new labor mix be continued?

3.

Compute the variable overhead rate and efficiency variances.
Answer #1

1.Material Price Variance = (Standard Price – Actual Price)*Actual Quantity Purchased

= (25-Actual Price)*12,000

= 25*12,000 – 282,000

= $18,000 F

Material Quantity Variance = (Standard Quantity – Actual Quantity Used)*Standard Price

= (4,100*2.20 –9,250)*25

= $5,750 U

Yes, since overall positive material variance

2.a. Labor Rate variance = (Standard Rate – Actual Rate)*Actual Hours

= (15-11.50)*110*25

= $9,625 F

Labor Efficiency Variance = (Standard Hours – Actual Hours)*Standard Rate

= (4,100*0.5 – 2,750)*15

= $10,500 U

No, should not be continued as it lead to drastic fall in efficiency

3-a Overhead Rate Variance = (Standard Rate – Actual Rate)*Actual Hours

= (3-Actual Rate)*2750

= $5850 F

Overhead Efficiency Variance = (Standard Hours – Actual Hours)*Standard Rate

= (-700)*3

= $2,100 U